You’ve observed the following returns on Crash-n-Burn Computer’s stock over the past five years: 4 percent, -12 percent, 25 percent, 21 percent, and 17 percent.

  

Requirement 1:
What was the arithmetic average return on Crash-n-Burn’s stock over this five year period?
(Click to select)  8.91%  13.09%  11.00%  11.99%  13.75%

   

Requirement 2:
(a)

What was the variance of Crash-n-Burn’s returns over this period? (Do not round intermediate calculations.)

(Click to select)  0.01843  0.01820  0.02370  0.02070  0.02275

   

(b)

What was the standard deviation of Crash-n-Burn’s returns over this period? (Do not round intermediate calculations.)

(Click to select)  17.95%  13.49%  12.22%  15.39%  15.08%
Answer #1

arithmetic average return =Total return/Total time period

=(4-12+25+21+17)/5=11%

Return (Return-Average Return)^2
4 (4-11)^2=49
-12 (-12-11)^2=529
25 (25-11)^2=196
21 (21-11)^2=100
17 (17-11)^2=36
Total=910%

Standard deviation=[Total (Return-Average Return)^2/(Time Period-1)]^(1/2)

=[910/(5-1)]^(1/2)

=[910/4]^(1/2)

=15.08%(Approx).

Variance=Standard deviation^2

=0.02275